OpenStudy (anonymous):

what is the solution of the following linear system? -4x-6y=8 -2x-3y=4

OpenStudy (anonymous):

-4x-6y=8 -6y=4x+8 -6y/6=-4x/6-8/6 y=2x/3-4/3 substitute, -2x-3(-2x/3-4/3)=4 -2x+6x/3+12/3=4 -2x+2x+4=4 -2x+2x=4-4 0=0 no solution